FEL Upgrade Project Weekly Brief – August 21–25, 2000


This week a number of the FEL team members presented key findings concerning the accelerator physics of the IR Demo FEL at the annual Linac Conference, in Monterey, CA.  Dave Douglas presented an invited talk on the recirculated, energy recovering linac in the IR Demo in addition to a poster on the design of the accelerator driver for the FEL upgrade. Presentations were also given by R. Li on our coherent synchrotron studies and Lia Merminga on rf stability.


 The monthly  report summarizing the technical, cost and schedule progress for the month of July for the IR Demo Upgrade was prepared and forwarded to ONR and DOE today.

WBS 3 (Beam Physics)

 The beam physics team was at the Linac conference this week as noted in the Highlights above.

WBS 4 (Injector)

We received a proposal from the William & Mary plasma group for ion implantation of the redesigned electrodes for the upgrade photoemission gun. The proposal is currently under review.  The test system for the apertured cesiator was successfully baked out without the film thickness monitor (which is in C. Sinclair’s hands for repair of the small leak in the feedthrough).  It looks like the cesiator qualifies for gun vacuum service.

WBS 6 (RF)

We have begun the installation of the low level rf racks in the zone 3 and 4 rf positions in the second floor gallery. The HPA rack is installed for zone 4.

WBS 8  (I & C)

 The weeks activities included significant progress in many areas. The new 3" Beam Position Monitor (BPM) has the initial welding complete as is being RF tested. RF tests are done prior to welding on the end nose pieces to ensure there is a good contact to the SMA RF feedthru and that there is an equal response between the four electrodes. The new 3" shielded beam viewer drawings are now complete and are being prepared as
packages to be sent out for quotes. The BPM drawings will be sent together with the viewers but as a separate package. At the FEL conference discussions continued on the new "analog" BPM electronics.  This is an active collaboration with the instrumentation group at the Rossendorf FEL facility. A first draft of these electronics will be
brought here and jointly tested during the October user run.

The final requirements for the DC power supplies was agreed upon and distributed. This will require 32 additional trim power supplies but only a single 5 kWatt larger supply needs to be purchased new. The balance will be from the currently used supplies and from power supplies that are available from other upgrade activities.  The number of shunts will raise to four from the current three and a pair of switches will be incorporated to switch between beam lines.

The Laser Safety System (LSS) smart card access system is being upgraded in preparation of bringing on the last two (of six) user labs. These upgrades will include the ability to change scan rates as the administrative level.  During the spring user run this system had crashed a number of times, in the ensuing period a card was attached to
a pneumatic cylinder and cycled through LabView 5000 times! There were NO failures. The system did however crash in the middle of the user run - it was traced to a user presenting his card like he was swatting at a
bee. We had slowed down the read times because of these crashes. Now we are speeding it back up having
retrained our users and labeled the readers for proper operation.  Additionally the LSS smoke detectors have been clearly labeled so there is no confusion.  These are interlocked to the laser to create a tiered approach to fire safety, they shut down only the photon beam in the event of excess smoke from an experiment.

The FEL Instrumentation & Control system documentation is being reviewed and "cleaned up". There have been a number of outstanding tasks that are being addressed as we gear up for the upgrade activities. A "short list"
has been chosen for the electronic technician position, candidates will be interviewed on Monday. Progress has been made on the list for the coordinator position, but no suitable candidates have been found for the
2 year term designer position.   Progress has also been made on cleaning up and organizing the FEL storage areas.

WBS 9 (Transport)
 We continued layout of the GX Prototype, nearly completing the final detail layout with such items as alignment cartridges, coil brackets, field clamp adjustment aids, brackets for incoming leads, thermal switch placement, circuitry and terminal strip. Analysis of 3D magnetic model of the GX yielded data indicating that, within the required 23 cm good field region, (The actual width is 19 cm, but our model is only good within 4 cm.), the core field is not the 1 part per 10,000 we require.  Since we already have included a large shelf on the square edged pole (the magnet is wide enough already!) to obtain a good field, we are calculating (getting promising results) the
effect of extending the field flatness width by using a simple bump on the edges of the pole. Quadrupoles  Four coils are wound and qualified.  After winding a fifth as a spare, the cooling plates will be added to complete the set.  The machining of the core prototype quadrant has been pulled from the shop and sent to an outside firm partially because of SNS interference.  However, this is wise because the outside firm has a milling machine that matches the computer controlled milling machine the shop is purchasing to complete the remaining 180 quadrupole quadrants.
Machining will start Sept. 5.  In the interim, the outside shop has machined a trial piece of just the pole face, which is being evaluated by the survey group.


Butch Dillon-Townes started a Watson Chart for tracking the progress on mechanical hardware that is resident in the shared folder for the group.

Completion of the report of the Technical Evaluation Team for the engineering services contract was delayed by the FEL Conference and remains nearly complete. Overall layouts continued to concentrate on integrating the BPMs, the quadrupoles, the existing inventory of corrector dipoles and the beam viewers in to a compact girder unit.

 WBS 10 (Wiggler)

Hoses, thermal switches and bus bar have been installed on the Northrop

Grumman Wiggler and manifold fabrication has started.  The drawings for the dispersion section have been sent out for a budgetary price.

WBS 11 (Optics)

Received fabricated adapter to attach a pyroelectric array to the spectrometer. Confirmed adapter fit.
Worked further with engineering and optics group staff to develop the FEL cavity mirror test stand.
Held a brainstorming session of the optics group to share information gleaned from recent trips to TRW, LLNL, and the FEL Conference/Workshop.


The only operations this week was a planned diagnostic test for the intermittent arcing of the HVPS for the gun.  After finding colorless deposits on the fiberglass support rods within the coaxial elbow which connects the HVPS to the gun, these rods were replaced. In subsequent tests of the gun on Friday afternoon, there was no sign of arcing at the nominal 320 kV operating voltage.

Privacy and Security Notice